# A Novel Lexicon for the Moral Foundation of Liberty

### v1

@@ Abstract
The moral value of liberty is a central concept in our inference system when it comes to taking a stance towards controversial social issues such as vaccine hesitancy, climate change, or the right to abortion. Here, we propose a novel Liberty lexicon evaluated on more than 3,000 manually annotated data both in in- and out-of-domain scenarios. As a result of this evaluation, we produce a combined lexicon that constitutes the main outcome of this work. This final lexicon incorporates information from an ensemble of lexicons that have been generated using word embedding similarity (WE) and compositional semantics (CS). Our key contributions include enriching the liberty annotations, developing a robust liberty lexicon for broader application, and revealing the complexity of expressions related to liberty across different platforms. Through the evaluation, we show that the difficulty of the task calls for designing approaches that combine knowledge, in an effort of improving the representations of learning systems.
Keywords lexical resources, moral foundation theory, natural language processing
@@ 1 Introduction
Moral values are fundamental to our decision-making process, especially regarding controversial social issues. When taking a stance, for instance, on global warming or vaccine adherence, we consult - consciously or unconsciously - our moral system of values. The Moral Foundations Theory (MFT) was created precisely to explain morality across cultures Haidt and Joseph (2004), proposing five foundations, namely care, fairness, loyalty, authority and sanctity. In a much later revision, the theory was enhanced with a new sixth dimension: liberty (Haidt, 2012). The “Liberty/Oppression” foundation is about people’s reactance and resentment towards those who dominate them and restrict their liberty. Moral notions captured by the Liberty foundation include freedom of choice and individual responsibility of actions, which repeatedly emerged as fundamental decision-making drivers of crucial prosocial behaviors such as vaccine adherence (Amin et al., 2017; Beiró et al., 2023; Zhang et al., 2023), and cooperation during crisis (Mejova et al., 2023).
Recent works focus on the automatic detection of moral values in text, employing annotated lexicons, either for unsupervised detection or as features in a learning system (Mooijman et al., 2018; Rezapour et al., 2021; Kennedy et al., 2021; Preniqi et al., 2021; Mejova et al., 2023; Mokhberian et al., 2020; Zhang et al., 2023). Given that the liberty foundation was added to the MFT theory subsequently, there were initially no linguistic resources for it. Preliminary approaches to liberty assessment were based on purely data-driven lexical characterization Araque et al. (2021, 2022), lacking, however, a solid evaluation against a benchmarked ground-truth. This is precisely the gap we are addressing in this study.
@@
We gathered data from various platforms to cover multiple aspects of the expression of the liberty foundation. In particular, we included (i) the Wikipedia [^1] https://www.wikipedia.org and Conservapedia [^2] https://www.conservapedia.com projects, encyclopedia projects of general content but diverse viewpoints, (ii) the r/Libertarian and r/Conservative communities on Reddit.com, forums of political discussion of general interest, (iii) the Black Lives Matter (BLM) and Election (Elect.) datasets from the MFTC Twitter Corpus Hoover et al. (2020), a collection of tweets discussing racial discrimination and the US Presidential Elections of 2016, respectively, as well as (iv) posts and comments from META’s Pages regarding the vaccination debate (Vaccine). The first two scenarios (Wikipedia vs Conservapedia and Libertarian vs Conservative on Reddit) act as “natural experiments” expressing the viewpoints of communities with diverse opinions and stances on the liberty foundation described by the MFT framework. To ensure a robust ground-truth, we obtained manual annotations of the expression of the moral foundation of liberty in the BLM, Elect., and META’s vaccination-related posts and comments.
@@
We generated two lexicons per dataset, employing two complementary approaches; the word embedding similarity (WE) (Turney and Littman, 2003) and the compositional semantics (CS) (Liang et al., 2013). The first automatically extracts a set of seed words using frequency shifts, comparing new words’ embeddings to seed words’ embeddings to determine their alignment with the foundation’s principles. The second method assumes that each word expresses the side of the foundation more frequently present in the documents where the word appears. We explore a lexicon aggregation approach based on overlapping terms that combines the benefits of the two methods. We also propose a combined representation approach which takes into consideration the individual lexical resources while accounting for overfitting issues. Finally, we evaluate the lexicons obtained per dataset, both in cross-domain experimental setups and out-of-domain ones.
We contribute to the state of the art in moral foundation recognition in the following ways. We expand the benchmark dataset of the Moral Foundations Theory Corpus (MFTC) by providing valuable manual annotations on the liberty moral foundation and rendering them available to the scientific community. Our work developed a refined and versatile liberty lexicon capable of effectively generalizing over previously unseen domains; the final liberty lexicon as well as the intermediate ones will also be available online. Our research sheds light on the nuanced variations in the expression of liberty across different domains, providing valuable insights into how this critical moral foundation can manifest differently within diverse contexts. These contributions collectively enhance our understanding of moral analysis and pave the way for more accurate and comprehensive evaluations.
@@ 2 Related Works
The Moral Foundations Dictionary (MFD) Graham et al. (2009) is a collection of lemmas and associated moral traits, assembled by experts and typically used together with the Linguistic Inquiry and Word Count (LIWC) software Tausczik and Pennebaker (2010) to estimate moral traits and investigate differences in moral concerns between different cultural groups. Garten et al. (2018) proposed the Distributed Dictionary Representations (DDR) method based on psychological dictionaries and semantic similarity to quantify the presence of moral sentiment around a given topic. Later on, the authors extended the method, incorporating demographic embeddings into the language representations (Garten et al., 2019).
@@
In an attempt to address several of the limitations of the MFD, Araque et al. (2020) proposed a data-driven generated lexicon, the MoralStrength, which expanded the original MFD employing the WordNet synsets and crowdsourced annotations. Different from the MFD, where each foundation is considered a bipolar of “virtue” and “vice”, MoralStrength treats each foundation as a continuum, assigning a numeric value of moral valence to each lemma that indicates the weight with which the lemma is expressing the specific value. Hopp et al. (2021) developed the extended Moral Foundations Dictionary (eMFD), a lexicon which expands the MFD based on crowdsourced annotations. Each lemma in eMFD is assigned a continuously weighted vector that expresses the probability that the lemma belongs to any of the five moral foundations.
@@
Notably, none of the above lexicons though included the liberty moral foundation. A first attempt to derive a lexicon from assessing the presence of liberty in the text was presented by Araque et al. (2021). They considered pairs of Wikipedia Pages and their Conservapedia counterparts as natural expressions of the liberty-oppression divide. They created a series of word embeddings which were then compared through cosine similarity to a set of seed words defined by experts to generate a lexicon. Their design comes with the obvious conceptual limitations of considering the Wikipedia project as expressing a strongly libertarian position and initiating the embeddings with a list of manually selected seed words from expert annotators. More recently,  Araque et al. (2022) proposed a liberty lexicon generation approach based aligning documents from online news sources with different worldviews. The LibertyMFD was later employed by Araque et al. (2023) to fine-tune the approach proposed by Consoli et al. (2022) for analysing how the Spanish news cover the female (un)employment topic in terms of sentiment and moral values, as well as how this sentiment evolves over time.
@@
Although pioneering, their approach suffered from the lack of a solid ground-truth on which to evaluate the generated lexicons. Due to the lack of a ground-truth dataset, the lexicon evaluation was based on the assumption that news from different political orientations would express opposite notions with respect to the liberty moral foundation. Instead, here, we evaluate the lexicons against solid manual annotations for the liberty foundation of the benchmark Moral Foundations Twitter Corpus datasets (BLM and Elect.) which we render available to the scientific community.
@@ 3 Data Collection
Moral Foundations Twitter Corpus (MFTC)
The Moral Foundations Twitter Corpus is a corpus consisting of seven independent datasets (35k tweets in total), manually annotated for the original five moral foundations Hoover et al. (2020), but not the liberty foundation.
The Black Lives Matter Twitter Corpus (BLM) van der Veen (2022) and the Elections Corpus (Elect.) Davidson et al. (2017) are the two largest datasets in this collection, and we manually annotated them [^3] We did not proceed to the annotation of the entire MFTC corpus due to funding limitations. as per the moral foundation of Liberty, relying on a popular tool for crowdsourcing and human validation, i.e., Amazon Mechanical Turk (AMT hereafter) provided by the Amazon SageMaker Ground Truth service. The Black Lives Matter Twitter Corpus focuses on tweets specifically regarding the Black Lives Matter movement, and it contains 4,352 tweets. The Election corpus relates to the US 2016 Presidential election, and consists of 4,370 tweets.
To ensure coherence, we followed the same procedure and annotation scheme that Hoover et al. (2017) used for the MFTC annotation. Moreover, inspired by the annotation approach of the MoralStrength lexicon Araque et al. (2020), we added the notion of “strength”, which indicates the degree to which each lemma expresses the liberty moral foundation in addition to its presence and polarity.
We assigned each tweet to nine independent annotators and asked them to rate the extent to which each tweet expressed a “Liberal” or “Oppressive” moral value on a scale from 1 to 9. The score magnitude represents the intensity of the Liberty/Oppression expressed in a tweet, as perceived by the annotator: a score close to 9 indicates that the sentence expresses a highly oppressive connotation, while a score value close to 1 is associated with a very libertarian connotation. Should the sentence not be associated with neither an oppressive nor a libertarian connotation, then the annotator could assign a neutral score. The intercoder agreement score provided by AMT is 92% [^4] Annotations will be available upon acceptance..
@@
Reddit is increasingly becoming a reliable data source in computational studies (Proferes et al., 2021). Aiming to profile the language of libertarian and conservative users, we extracted textual content from the r/Libertarian and r/Conservative communities, which are self-proclaimed networks of libertarian and conservative ideas, respectively. Initially, we considered posts and comments published between August 2008 and April 2021, obtaining overall 1,127,005 documents. From these, we have filtered empty and other unusable content, and undersampled the rest to obtain a final amount of 100,000 instances.
@@
We use the dataset described in Araque et al. (2021), based on page alignment between Wikipedia and Conservapedia according to their title (henceforth the WikiCon dataset). More than 37,000 articles between Wikipedia and Conservapedia have been aligned, of which approximately 28,000 pages had identical titles, and the remaining were aligned based on redirect pages. The entire corpus contains 106 million tokens and 558,000 unique words. The dataset has been filtered using page categories related to politics, while a length ratio filter has been applied between Wikipedia and Conservapedia documents to improve dataset quality. This ratio compares the number of words in a Wikipedia document to the number of terms in the corresponding Conservapedia document, and excludes the pairs with ratio higher than 10, resulting in 57,078 documents split equally between 28,539 Conservapedia and Wikipedia sources.
@@
Finally, we use a dataset on vaccinations, which comprises anonymous posts and comments from about 200 Facebook Pages, collected through the Facebook API from January 2012 to June 2019 Prado et al. (2022). The total number of comments and posts from both sides of the vaccine debate amounts to 607,105. The creators of the dataset randomly selected approximately 1,500 comments and manually annotated the presence of the liberty moral foundation in the snippet, indicating also the polarity of the foundation as “virtue” (liberty) or “vice” (oppression). A summary of the datasets used in this work is presented in Table 1.
@@ 4.2 Lexicon Generation
Word Embedding Similarity
Based on the approach proposed by Turney and Littman (2003), our first strategy for generating lexicons relies on word embedding similarity between the vectors of the positive and negative instances of a dataset’s documents. Hence, the method relies on a set of seed words that accurately represent the domains we aim to differentiate. However, arbitrary selection of seed words can bias the output, since variations in the seed word list lead to differences in the final lexicon. To overcome this issue, we obtained the set of seed words in a data-driven way by estimating the frequency shifts (Gallagher et al., 2021) of the lemmas between the positive and negative documents, as done in Araque et al. (2022). This approach helps us to avoid the limitations of arbitrarily selecting the seed words. Thus, we consider the relative frequency of a word $w$ in a set of documents $D$:
@@
The seed word lists are generated based on prominent differences in word frequency shifts. We apply a minimum frequency threshold of 100 to filter out less common lemmas. We then use the word2vec algorithm (Le and Mikolov, 2014) to compute the vector for each word, using the standard parameter setting and a vector dimension of 300. The lexicon is generated by estimating the cosine similarity between the word vectors obtained using the emerging seed words. To compute the moral polarity of a word $w_{i}$ from the documents, we use the sets of seed words for the “oppressive” orientation ($S_{C}$) and the “liberty” direction ($S_{L}$), and estimate the polarity based on the cosine similarity:
@@
where sim represents the cosine similarity as estimated by the word embedding model. The obtained polarity is positive if $w_{i}$ is related to the positive seed words and a negative value if the word is more related towards the negative seed words. For the rest of the paper, we refer to this model as the WE model and we generate one lexicon for each dataset (except Vaccine, which is only used for testing).
Compositional Semantics
The second approach involves using the Compositional Semantics (CS) method (Liang et al., 2013), previously used to generate emotion lexicons Staiano and Guerini (2014); Araque et al. (2019). The CS method applies a projection of moral values from a document to its words. The underlying assumption is that each word is associated with the moral value present in the documents where the word appears more frequently.
To generate a word-by-moral association matrix ($M_{WM}$), we first create a document-by-moral matrix $M_{DM}$, which shows the distribution of the liberty foundation across the training dataset. We then generate a word-by-document matrix $M_{WD}$, which indicates the number of occurrences for each word in the vocabulary within a given document, normalized by the total number of words per document. To obtain the word-by-moral matrix, we perform a multiplication using the following expression:
@@
Using this approach, words and their corresponding value of liberty can be merged by calculating the product of the weight of a word and the weight of the moral value in each document. The resulting scores are then normalised (column-wise), over-representation issues are addressed, and each lemma is scaled (row-wise) to sum up to one. Previous validation of lexicons has shown that this normalisation approach is suitable Araque et al. (2019, 2022). This approach is referred to as the CS model and we generate one lexicon for each dataset except Vaccine.
Overlap Lexicon
The domain-specific lexicons express the liberty dimension that is dependent on the topics of the dataset on which they are generated from. However, we are interested in deriving a general, higher-level representation, so that the final users of the resource have a unified and domain-independent resource. To this regard, we synthesise a unified resource merging the obtained discrete lexicons. This approach (i) augments the coverage of the consolidated lexicon and (ii) discards uncommon tokens and their annotations. The basic process to obtain such a lexicon starts by defining a unified vocabulary as the union of the vocabularies of all individual lexicons. This union can be controlled with a selection parameter expressed as a percentage value. That is, if we define a selection parameter of 50%, a word would be included in the union of vocabularies if it appears in at least the 50% of all considered lexicons.
Then, we align the numeric assignment each token has in the individual lexicon. We estimate the average score of these assignments, incorporating them into the unified resource, if the volume of annotations satisfies the threshold stipulated by the chosen proportion (selection parameter). We denote the obtained lexicon as Lexicon Overlap (avg.) [^5] The lexicon will be released upon acceptance..

### v2

# LML: A Novel Lexicon for the Moral Foundation of Liberty

@@ Abstract
The moral value of liberty is a central concept in our inference system when it comes to taking a stance towards controversial social issues such as vaccine hesitancy, climate change, or the right to abortion. Here, we propose a novel Liberty lexicon evaluated on more than 3,000 manually annotated data both in in- and out-of-domain scenarios. As a result of this evaluation, we produce a combined lexicon, referred to as the Liberty Moral Lexicon (LML), that constitutes the main outcome of this work. LML incorporates information from an ensemble of lexicons that have been generated using word embedding similarity (WE) and compositional semantics (CS). Our key contributions include enriching the liberty annotations, developing a robust liberty lexicon for broader application, and revealing the complexity of expressions related to liberty across different platforms. Through the evaluation, we show that the difficulty of the task calls for designing approaches that combine knowledge, in an effort of improving the representations of learning systems.
Keywords:
Natural language, moral foundations, liberty, oppression.
@@ 1 Introduction
Moral values are fundamental to our decision-making process, especially regarding controversial social issues. When taking a stance, for instance, on global warming or vaccine adherence, we consult - consciously or unconsciously - our moral system of values. The Moral Foundations Theory (MFT) was created precisely to explain morality across cultures [16], proposing five foundations, namely care, fairness, loyalty, authority and sanctity. In a much later revision, the theory was enhanced with a new sixth dimension: liberty [15]. The “Liberty/Oppression” foundation is about people’s reactance and resentment towards those who dominate them and restrict their liberty. Moral notions captured by the Liberty foundation include freedom of choice and individual responsibility of actions, which repeatedly emerged as fundamental decision-making drivers of crucial prosocial behaviors such as vaccine adherence [2, 8, 33], and cooperation during crisis [26]. Recent works focus on the automatic detection of moral values in text, employing annotated lexicons, either for unsupervised detection or as features in a learning system [27, 30, 22, 28, 26, 33]. Given that the liberty foundation was added to the MFT theory subsequently, there were initially no linguistic resources for it. Preliminary approaches to liberty assessment were based on purely data-driven lexical characterization [6, 7], lacking, however, a solid evaluation against a benchmarked ground-truth. This is precisely the gap we are addressing in this study.
@@
We gathered data from various platforms to cover multiple aspects of the expression of the liberty foundation. In particular, we included (i) the Wikipedia [^1] https://www.wikipedia.org and Conservapedia [^2] https://www.conservapedia.com projects, encyclopedia projects of general content but diverse viewpoints, (ii) the r/Libertarian and r/Conservative communities on Reddit.com, forums of political discussion of general interest, (iii) the Black Lives Matter (BLM) and Election (Elect.) datasets from the MFTC Twitter Corpus [19], a collection of tweets discussing racial discrimination and the US Presidential Elections of 2016, respectively, as well as (iv) posts and comments from META’s Pages regarding the vaccination debate (Vaccine). The first two scenarios (Wikipedia vs Conservapedia and Libertarian vs Conservative on Reddit) act as “natural experiments” expressing the viewpoints of communities with diverse opinions and stances on the liberty foundation described by the MFT framework. To ensure a robust ground-truth, we obtained manual annotations of the expression of the moral foundation of liberty in the BLM, Elect., and META’s vaccination-related posts and comments.
@@
We generated two lexicons per dataset, employing two complementary approaches; the word embedding similarity (WE) [32] and the compositional semantics (CS) [24]. The first automatically extracts a set of seed words using frequency shifts, comparing new words’ embeddings to seed words’ embeddings to determine their alignment with the foundation’s principles. The second method assumes that each word expresses the side of the foundation more frequently present in the documents where the word appears. We explore a lexicon aggregation approach based on overlapping terms that combines the benefits of the two methods, resulting in the Liberty Moral Lexicon (LML). We also propose a combined representation approach which takes into consideration the individual lexical resources while accounting for overfitting issues. Finally, we evaluate the lexicons obtained per dataset, both in cross-domain experimental setups and out-of-domain ones.
We contribute to the state of the art in moral understanding by augmenting the benchmark dataset of the Moral Foundations Theory Corpus (MFTC) with valuable manual annotations on the liberty moral foundation and rendering them available to the scientific community. Our novel Liberty Moral Lexicon combines a variety of data sources covering a wide range of topics, resulting into a refined and versatile liberty lexicon capable of effectively generalizing over previously unseen domains [^3] The LML dictionary is available at the following link: https://github.com/oaraque/moral-foundations/tree/master/liberty/3rd˙version. Moreover, LML enables researchers to bridge qualitative and quantitative methods, enhancing tools like TOMEA [25] in uncovering the contextual variability of moral language. These contributions collectively enhance our understanding of moral analysis and pave the way for more accurate and comprehensive evaluations.
@@ 2 Related Works
The Moral Foundations Dictionary (MFD) [14] is a collection of lemmas and associated moral traits, assembled by experts and typically used together with the Linguistic Inquiry and Word Count (LIWC) software [31] to estimate moral traits and investigate differences in moral concerns between different cultural groups. Garten et al. [12] proposed the Distributed Dictionary Representations (DDR) method based on psychological dictionaries and semantic similarity to quantify the presence of moral sentiment around a given topic. Later on, the authors extended the method, incorporating demographic embeddings into the language representations [13].
@@
In an attempt to address several of the limitations of the MFD, [5] proposed a data-driven generated lexicon, the MoralStrength, which expanded the original MFD employing the WordNet synsets and crowdsourced annotations. Different from the MFD, where each foundation is considered a bipolar of “virtue” and “vice”, MoralStrength treats each foundation as a continuum, assigning a numeric value of moral valence to each lemma that indicates the weight with which the lemma is expressing the specific value. [21] developed the extended Moral Foundations Dictionary (eMFD), a lexicon which expands the MFD based on crowdsourced annotations. Each lemma in eMFD is assigned a continuously weighted vector that expresses the probability that the lemma belongs to any of the five moral foundations.
@@
Notably, none of the above lexicons though included the liberty moral foundation. A first attempt to derive a lexicon from assessing the presence of liberty in the text was presented by [6]. They considered pairs of Wikipedia Pages and their Conservapedia counterparts as natural expressions of the liberty-oppression divide. They created a series of word embeddings which were then compared through cosine similarity to a set of seed words defined by experts to generate a lexicon. Their design comes with the obvious conceptual limitations of considering the Wikipedia project as expressing a strongly libertarian position and initiating the embeddings with a list of manually selected seed words from expert annotators. More recently,  [7] proposed a liberty lexicon generation approach based aligning documents from online news sources with different worldviews. The LibertyMFD was later employed in [4] to fine-tune the approach proposed by Consoli et al. [9] for analysing how the Spanish news cover the female (un)employment topic in terms of sentiment and moral values, as well as how this sentiment evolves over time.
@@
Although pioneering, their approach suffered from the lack of a solid ground-truth on which to evaluate the generated lexicons. Due to the lack of a ground-truth dataset, the lexicon evaluation was based on the assumption that news from different political orientations would express opposite notions with respect to the liberty moral foundation. Instead, in LML, we evaluate the lexicons against solid manual annotations for the liberty foundation of the benchmark Moral Foundations Twitter Corpus datasets (BLM and Elect.) which we render available to the scientific community.
@@ 3 Data Collection
Moral Foundations Twitter Corpus (MFTC).
The MFTC is a corpus consisting of seven independent datasets (35k tweets in total), manually annotated for the original five moral foundations [19], but not the liberty foundation. The Black Lives Matter Twitter Corpus (BLM) and the Elections Corpus (Elect.) are the two largest datasets in this collection, and we manually annotated them [^4] We did not proceed to the annotation of the entire MFTC corpus due to funding limitations. as per the moral foundation of Liberty, relying on a popular tool for crowdsourcing and human validation, i.e., Amazon Mechanical Turk (AMT hereafter) provided by the Amazon SageMaker Ground Truth service. The Black Lives Matter Twitter Corpus focuses on tweets specifically regarding the Black Lives Matter movement, and it contains 4,352 tweets. The Election corpus relates to the US 2016 Presidential election, and consists of 4,370 tweets. To ensure coherence, we followed the same procedure and annotation scheme of Hoover et al. [20] used for the MFTC annotation. Moreover, inspired by the annotation approach of the MoralStrength lexicon [5], we added the notion of “strength”, which indicates the degree to which each lemma expresses the liberty moral foundation in addition to its presence and polarity. We assigned each tweet to nine independent annotators and asked them to rate the extent to which each tweet expressed a “Liberal” or “Oppressive” moral value on a scale from 1 to 9. The score magnitude represents the intensity of the Liberty/Oppression expressed in a tweet, as perceived by the annotator: a score close to 9 indicates that the sentence expresses a highly oppressive connotation, while a score value close to 1 is associated with a very libertarian connotation. Should the sentence not be associated with neither an oppressive nor a libertarian connotation, then the annotator could assign a neutral score. The intercoder agreement score provided by AMT is 92% [^5] Annotations will be available upon acceptance..
@@
Reddit is increasingly becoming a reliable data source in computational studies [29]. Aiming to profile the language of libertarian and conservative users, we extracted textual content from the r/Libertarian and r/Conservative communities, which are self-proclaimed networks of libertarian and conservative ideas, respectively. Initially, we considered posts and comments published between August 2008 and April 2021, obtaining overall 1,127,005 documents. From these, we have filtered empty and other unusable content, and undersampled the rest to obtain a final amount of 100,000 instances.
@@
We use the dataset described in [6], based on page alignment between Wikipedia and Conservapedia according to their title (henceforth the WikiCon dataset). More than 37,000 articles between Wikipedia and Conservapedia have been aligned, of which approximately 28,000 pages had identical titles, and the remaining were aligned based on redirect pages. The entire corpus contains 106 million tokens and 558,000 unique words. The dataset has been filtered using page categories related to politics, while a length ratio filter has been applied between Wikipedia and Conservapedia documents to improve dataset quality. This ratio compares the number of words in a Wikipedia document to the number of terms in the corresponding Conservapedia document, and excludes the pairs with ratio higher than 10, resulting in 57,078 documents split equally between 28,539 Conservapedia and Wikipedia sources.
@@
Finally, we use a dataset on vaccinations, which comprises anonymous posts and comments from about 200 Facebook Pages, collected through the Facebook API from January 2012 to June 2019 [8]. The total number of comments and posts from both sides of the vaccine debate amounts to 607,105. The creators of the dataset randomly selected approximately 1,500 comments and manually annotated the presence of the liberty moral foundation in the snippet, indicating also the polarity of the foundation as “virtue” (liberty) or “vice” (oppression). A summary of the datasets used in this work is in Table 1.
@@ 4.2 Lexicon Generation
Word Embedding Similarity.
Based on the approach proposed in [32], our first strategy for generating lexicons relies on word embedding similarity between the vectors of the positive and negative instances of a dataset’s documents. Hence, the method relies on a set of seed words that accurately represent the domains we aim to differentiate. However, arbitrary selection of seed words can bias the output, since variations in the seed word list lead to differences in the final lexicon. To overcome this issue, we obtained the set of seed words in a data-driven way by estimating the frequency shifts [11] of the lemmas between the positive and negative documents, as done in [7]. This approach helps us to avoid the limitations of arbitrarily selecting the seed words. Thus, we consider the relative frequency of a word $w$ in a set of documents $D$:
@@
The seed word lists are generated based on prominent differences in word frequency shifts. We apply a minimum frequency threshold of 100 to filter out less common lemmas. We then use the word2vec algorithm [23] to compute the vector for each word, using the standard parameter setting and a vector dimension of 300. The lexicon is generated by estimating the cosine similarity between the word vectors obtained using the emerging seed words. To compute the moral polarity of a word $w_{i}$ from the documents, we use the sets of seed words for the “oppressive” orientation ($S_{C}$) and the “liberty” direction ($S_{L}$), and estimate the polarity based on the cosine similarity:
@@
where sim represents the cosine similarity as estimated by the word embedding model. The obtained polarity is positive if $w_{i}$ is related to the positive seed words and a negative value if the word is more related towards the negative seed words. We refer this model to as the WE model and generate one lexicon for each dataset (except Vaccine, which is only used for testing).
Compositional Semantics.
The second approach involves using the Compositional Semantics (CS) method [24], previously used to generate emotion lexicons [3]. The CS method applies a projection of moral values from a document to its words. The underlying assumption is that each word is associated with the moral value present in the documents where the word appears more frequently. To generate a word-by-moral association matrix ($M_{WM}$), we first create a document-by-moral matrix $M_{DM}$, which shows the distribution of the liberty foundation across the training dataset. We then generate a word-by-document matrix $M_{WD}$, which indicates the number of occurrences for each word in the vocabulary within a given document, normalized by the total number of words per document. To obtain the word-by-moral matrix, we perform a multiplication using the following expression:
@@
Using this approach, words and their corresponding value of liberty can be merged by calculating the product of the weight of a word and the weight of the moral value in each document. The resulting scores are then normalised (column-wise), over-representation issues are addressed, and each lemma is scaled (row-wise) to sum up to one. Previous validation of lexicons has shown that this normalisation approach is suitable [3, 7]. This approach is referred to as the CS model and we generate one lexicon for each dataset except Vaccine.
Liberty Moral Lexicon (LML).
The domain-specific lexicons capture topic-dependent expressions of the liberty dimension based on their source datasets. However, we are interested in deriving a general, higher-level representation, so that the final users of the resource have a unified and domain-independent resource. For this purpose, we synthesise a unified resource merging the obtained discrete lexicons, obtaining our Liberty Moral Lexicon. This approach (i) augments the coverage of the consolidated lexicon and (ii) discards uncommon tokens and their annotations. The basic process to obtain such a lexicon starts by defining a unified vocabulary as the union of the vocabularies of all individual lexicons. This union can be controlled with a selection parameter expressed as a percentage value. That is, if we define a selection parameter of mm%, a word would be included in the union of vocabularies if it appears in at least the mm% of all considered lexicons. Then, we align the numeric assignment each token has in the individual lexicon. We estimate the average score of these assignments, incorporating them into the unified LML resource [^6] The Liberty Moral Lexicon (LML) will be publically released upon acceptance., if the volume of annotations satisfies the threshold stipulated by the chosen proportion (selection parameter).